# Reply to Comment on “Properties and dynamics of generalized squeezed states”

### v1

@@ Abstract
In our paper Ashhab2025 , our numerical simulations showed that, unlike displacement and conventional squeezing, higher-order squeezing exhibits oscillatory dynamics. Subsequently, Gordillo and Puebla pointed out that simulation results depend on whether the state space in the simulations is even or odd Gordillo . Using additional derivations, they argued that the oscillatory dynamics is unphysical and that the photon number must increase monotonically as a function of the squeezing parameter $r$. We agree with the observation of an even-odd parity dependence in the simulations. We independently noticed the same feature in our simulations after the publication of Ref. Ashhab2025 . This observation led us to perform a more detailed investigation of the numerical simulation and mathematical aspects of the generalized squeezing problem. Our new findings were reported in Ref. Ashhab2026 . Further analysis was reported in Ref. Fischer . Our conclusion is that the generalized squeezing operator is physically not well defined but can be made well defined when combined with additional information about the physical system under study. We demonstrated this point in the case where we include an additional nonlinear interaction term in the Hamiltonian. We disagree with the claim that the photon number must be a monotonically increasing function of $r$. This claim contradicts the mathematically rigorous results of Ref. Fischer . Furthermore, we show that the oscillatory behaviour persists in two closely related, well-behaved models.
@@
In mathematical terms, the squeezing Hamiltonian is not essentially self-adjoint on the finite photon space Ashhab2026 ; Fischer . The fact that various quantities converge to well-defined asymptotic values for the even-size or odd-size simulations indicates that these two case (even and odd sizes) correspond to two different self-adjoint extensions of the Hamiltonian.
@@
The authors of the Comment Gordillo performed a calculation of the average photon number using a power series expansion of an operator combination that contains the squeezing operator. Using this calculation, they concluded that the number of photons must increase monotonically as a function of the squeezing parameter. This calculation is interesting and might contain useful information. However, as was pointed out in Ref. Fisher and reiterated in Refs. Ashhab2026 ; Fischer , using standard power series expansions when dealing with the problem of generalized squeezing can lead to non-convergent series. The generating Hamiltonian for the generalized squeezing operator, $i\left[\left(\hat{a}^{\dagger}\right)^{n}-\hat{a}^{n}\right]$, is not self-adjoint on the space spanned by the untruncated Fock state basis when $n\geq 3$. In other words, the Hamiltonian is physically not well defined on the physically intuitive state space spanned by photon number states ranging from zero to infinity. To make the squeezing operator unitary and physically meaningful, one must deal with subtle mathematical questions to find a domain (or state space) on which the Hamiltonian is self-adjoint. Similarly, one must be careful when performing power series expansions for this model, and one must in general provide convergence proofs to support the correctness of the conclusions. Any results obtained using non-convergent series expansions cannot be considered reliable. Our results in Ref. Ashhab2026 show that the regularized dynamics that include a Kerr term are convergent and consistently exhibit oscillatory behaviour. Furthermore, the rigorous analysis in Ref. Fischer leads to the conclusion that the photon number cannot increase monotonically. Since the conclusion of the Comment Gordillo , which is based on a heuristic series expansion, contradicts our mathematical analysis and the simulation-size-independent results of our recent simulations, we believe that the conclusion of the Comment is incorrect.
@@
One important advantage of using the quantum-pump model is that it does not have any serious issues with self-adjointness, and hence we do not need to deal with subtle mathematical questions such as finding a self-adjoint extension defined on a carefully chosen domain Hillery . We therefore eliminate a serious mathematical problem that plagues the generalized squeezing operator in the classical-pump model. Furthermore, this approach eliminates the need to choose a simulation truncation size arbitrarily. The number of photons in the quantum pump field introduces a natural cutoff: if the pump field runs out of photons, no more signal photons can be created. More specifically, the dynamics will be restricted to the state space $\left\{\left|0,N\right\rangle,\left|n,N-1\right\rangle,\left|2n,N-2\right\rangle,\cdots,\left|nN,0\right\rangle\right\}$. Any simulation that contains all of these states will give the same results. Adding more quantum states with higher photon numbers in the simulation will not affect the results at all.
We can now ask what phenomena are obtained in this well-behaved model. These phenomena include the oscillatory dynamics, as well as the strong parity dependence of the dynamics. We have verified with numerical simulations that both of these phenomena appear in the quantum-pump model. The parity dependence is particularly remarkable in this context. No matter how large the initial number of photons in the pump field is, the ensuing dynamics depends strongly on whether this number is even or odd. It should be noted here that even in the case $n=3$, the average number of signal photons scales roughly as $N^{0.5}$ (since the scaling should be similar to that in Fig. 8(a) in our paper Ashhab2025 , even though the truncation there is done arbitrarily), such that the pump field is hardly depleted on average. Nevertheless, the dynamics is sensitive to the parity of the initial number.

### v2

@@ Abstract
In our paper Ashhab2025 , our numerical simulations showed that, unlike displacement and conventional squeezing, higher-order squeezing exhibits oscillatory dynamics. Subsequently, Gordillo and Puebla pointed out that simulation results depend on whether the size of the state space in the simulations is even or odd Gordillo . Using additional derivations, they argued that the oscillatory dynamics is unphysical and that the photon number must increase monotonically as a function of the squeezing parameter $r$. We agree with the observation of an even-odd parity dependence in the simulations. We independently noticed the same feature in our simulations after the publication of Ref. Ashhab2025 . This observation led us to perform a more detailed investigation of the numerical simulation and mathematical aspects of the generalized squeezing problem. Our new findings were reported in Ref. Ashhab2026 . Further analysis was reported in Ref. Fischer . Our conclusion is that the generalized squeezing operator is physically not well defined but can be made well defined when combined with additional information about the physical system under study. We demonstrated this point in the case where we include an additional nonlinear interaction term in the Hamiltonian. We disagree with the claim that the photon number must be a monotonically increasing function of $r$. This claim contradicts the mathematically rigorous results of Ref. Fischer . Furthermore, we show that the oscillatory behaviour persists in two closely related, well-behaved models.
@@
In mathematical terms, the squeezing Hamiltonian is not essentially self-adjoint on the finite photon space Ashhab2026 ; Fischer . The fact that various quantities converge to well-defined asymptotic values for the even-size or odd-size simulations indicates that these two cases (even and odd sizes) correspond to two different self-adjoint extensions of the Hamiltonian.
@@
The authors of the Comment Gordillo performed a calculation of the average photon number using a power series expansion of an operator combination that contains the squeezing operator. Using this calculation, they concluded that the number of photons must be a monotonically increasing function of the squeezing parameter $r$. This calculation is interesting and might contain useful information. However, as was pointed out in Ref. Fisher and reiterated in Refs. Ashhab2026 ; Fischer , using standard power series expansions when dealing with the problem of generalized squeezing can lead to non-convergent series. The generating Hamiltonian for the generalized squeezing operator, $i\left[\left(\hat{a}^{\dagger}\right)^{n}-\hat{a}^{n}\right]$, is not self-adjoint on the space spanned by the untruncated Fock state basis when $n\geq 3$. In other words, the Hamiltonian is physically not well defined on the physically intuitive state space spanned by photon number states ranging from zero to infinity. To make the squeezing operator unitary and physically meaningful, one must deal with subtle mathematical questions to find a domain (or state space) on which the Hamiltonian is self-adjoint. Similarly, one must be careful when performing power series expansions for this model, and one must in general provide convergence proofs to support the correctness of the conclusions. Any results obtained using non-convergent series expansions cannot be considered reliable. Our results in Ref. Ashhab2026 show that the regularized dynamics that include a Kerr term are convergent and consistently exhibit oscillatory behaviour. Furthermore, our results in Ref. Ashhab2026 strongly suggest that the spectrum of the squeezing Hamiltonian is discrete. The discreteness of the spectrum was proved rigorously by the authors of Ref. Fischer . According to the quantum recurrence theorem, if the spectrum is discrete, no observable can be monotonic indefinitely. Since the conclusion of the Comment Gordillo , which is based on a heuristic series expansion, contradicts the results of rigorous mathematical analysis and the simulation-size-independent results of our recent simulations, we believe that the conclusion of the Comment is incorrect.
@@
One important advantage of using the quantum-pump model is that it does not have any serious issues with self-adjointness, and hence we do not need to deal with subtle mathematical questions such as finding a self-adjoint extension defined on a carefully chosen domain Hillery . We therefore eliminate a serious mathematical problem that plagues the generalized squeezing operator in the classical-pump model. Furthermore, this approach eliminates the need to choose a simulation truncation size arbitrarily. The initial number of photons in the quantum pump field introduces a natural cutoff: if the pump field runs out of photons, no more signal photons can be created. More specifically, the dynamics will be restricted to the state space $\left\{\left|0,N\right\rangle,\left|n,N-1\right\rangle,\left|2n,N-2\right\rangle,\cdots,\left|nN,0\right\rangle\right\}$. Any simulation that contains all of these states will give the same results. Adding more quantum states with higher photon numbers in the simulation will not affect the results at all.
We can now ask what phenomena are obtained in this well-behaved model. These phenomena include the oscillatory dynamics, as well as the strong parity dependence. We have verified with numerical simulations that both of these phenomena appear in the quantum-pump model. The parity dependence is particularly remarkable in this context. No matter how large the initial number of photons in the pump field is, the ensuing dynamics depends strongly on whether this number is even or odd. It should be noted here that even in the case $n=3$, the average number of signal photons scales roughly as $N^{0.5}$ (since the scaling should be similar to that in Fig. 8(a) in our paper Ashhab2025 , even though the truncation there is done arbitrarily), such that the pump field is hardly depleted on average. Nevertheless, the dynamics is sensitive to the parity of the initial number.